# Nonlocal Models in Biology and Life Sciences: Sources, Developments, and Applications

### v1

@@ 2.8 Nonlocal epidemiological models and nonlocality in immunology
Nonlocal epidemiological models have been used to study the spread of infectious diseases in biological systems, considering interactions beyond immediate neighbours and capturing the effects of long-range dispersal and spatial dependencies. In 1927, based on certain assumptions, Kermack and McKendrick proposed a deterministic SIR model to capture a transmitted viral or bacterial agent in a closed population kermack1927 ; ruan2007 . Here, we first describe the SIR model and its generalization, incorporating the space into the model. Let $S(t)$ be the density of a population susceptible to a disease but not yet infected at the time $t\geq 0$. Suppose $A(\theta)$ is the expected infectivity of an individual who has been infected $\theta$ time ago. Then, the Kermack-McKendrick model follows the integral differential equation kermack1927 :
@@
describes the epidemic’s growth in the initial phase of the disease propagation. The invasion criteria of the epidemic is $R_{0}$, which is greater than one. For the special choice $A(\theta)=\beta\exp(-\gamma\theta)$, where $\beta,\gamma>0$, the quantity
@@
$\frac{dR}{dt}$ $=\gamma I(t).$
In 1965, Kendall extended this SIR model (12) into a nonlocal epidemic model kendall1965 . Let $S(x,t)$, $I(x,t)$, and $R(x,t)$ be the densities of the susceptible, infected and removed individuals at the spatial location $x\in\mathbb{R}$ and time $t\geq 0$ such that their sum $S+I+R$ remains time-independent. We assume that the infected individuals are infectious and that the rate of infection is
@@
where $\beta$ is a positive constant and $\Psi$ is the normalized kernel function in $\mathbb{R}$. Taking these into the SIR model, the modified model becomes
@@
Nonlocal epidemiological models have been applied in biology to study the spread of infectious diseases in spatially extended and heterogeneous environments, such as urban areas with varying population densities or regions with different levels of healthcare access chen2021 . These models incorporate long-range interactions and spatial dependencies, allowing for a more realistic representation of disease transmission dynamics.
The COVID-19 pandemic was a global health crisis caused by the SARS-CoV-2 virus in a heterogeneous environment sohrabi2020 . This disease spreads rapidly through respiratory droplets when an infected person talks, coughs, or sneezes, and new variants of the virus continuously emerge, which causes concern among health experts around the world. Different types of measurements have been implemented to control the transmission of the virus, including targeted lockdowns, ramped-up testing, and widespread vaccination efforts. In terms of mathematical modelling, numerous mathematical frameworks have been proposed to investigate the transmission dynamics of the COVID-19 pandemic biswas2020 ; shi2023 ; xu2023 , and we highlight one of them where the spatial domain is considered. The novel coronavirus infects contacts during cold-chain transmission, so the study of the characteristics of COVID-19 dynamics can be more accurate in consideration of long-range population diffusion. Suppose $S(x,t)$, $V(x,t)$, $E(x,t)$, $I_{a}(x,t)$, and $I_{s}(x,t)$ are respectively for the populations of susceptible, fully vaccinated, latent infection, asymptomatic infected, and symptomatic infected at a point $x\in\mathbb{R}$ and time $t>0$, with $B(x,t)$ as the SARS-CoV-2 virus in the environment. Following shi2023 , we consider the spatio-temporal nonlocal mathematical model for COVID-19 as:
@@
where $\psi(x-y)$ is the probability that an individual at position $y$ moves to position $x$ in the space $\Omega$. The functions $f_{s}(x,t)$ and $f_{v}(x,t)$ are defined as
@@
where $\beta_{k}(k=e,a,s)$ are the rate of infection of infected people to susceptible people and $\alpha_{k}(k=e,a,s)$ are the rate of infection of infected people to those vaccinated, and $\beta_{b}$ and $\alpha_{b}$ are the rate of infection of the virus in the population. The other parameters used in the model (14) are given in shi2023 . The authors have shown that the effective rate of vaccination is one of the best measures to prevent and control the transmission of COVID-19 under current medical conditions.
Nonlocality in immunology refers to the phenomena in which immune system interactions and signalling processes extend beyond the immediate area of the immune cells involved. T-cell receptor (TCR) activation, for example, refers to the assumption that T-cell responses are not limited to the local contact between the TCR and its antigen. T-cell activation, on the other hand, is a networked and dynamic process that includes interactions at remote places smith2009 . It can lead to systemic effects, influencing immune responses and inflammation throughout the body, which has implications for autoimmune diseases and chronic inflammatory conditions bachmann2010 . Furthermore, nonlocal TCR activation is important in cancer immunotherapy because activated T cells may target cancer cells in distant metastatic locations. rosenberg2014 . In T memory cell homeostasis, for example, the bone marrow serves as a niche and interacts with stromal cells, and cytokines contribute to their survival and nonlocal homeostasis tokoyoda2009 . These T cells grow and contract in distinct stages, including interactions and signals that reach beyond the T cells’ local surroundings. During T cell activation and proliferation, cytokines such as interleukins are generated not just locally at the site of activation but also systemically. These cytokines assist in the nonlocal regulation of T-cell proliferation boyman2012 .
@@ 2.9 Relaxation-time models in life sciences
Relaxation-time models are pretty interesting, especially in the context of life sciences. These models help us understand the dynamics of biological systems and provide insights into the time-dependent behaviour of various processes. For instance, the nonlocal model (5) in space and time can be considered as the relaxation-time nonlocal model. A similar type of extension can be applied to the model (13). We first include the spatial variable into the model (13) to account for the random movement of individuals as de1979 :
@@
where the kernel function describes the interaction between the infective and the susceptible individuals at location $x$ and the present $t$, which occurred at location $y$ and at earlier instance $s$. This kernel function satisfies some conditions which are mentioned in ruan2007 .
The nonlocal effects and time delay may come in the model when the mature death and diffusion rates are age-independent. To derive such a delay model, we follow the work liang2003 where the author considered a single-species population model with two age classes and a fixed maturation period living in a spatial transport field. Suppose $u(x,a,t)$ is the population density at a spatial point $x\in\Omega\subset\mathbb{R}$ and time $t\geq 0$ with age $a\geq 0$. Following liang2003 , the single-species model given by
@@
where $v^{s}(x,t)=u(x,t-s,t)$.
@@ 2.10 Nonlocal problems and conservation laws
Nonlocal problems and conservation laws are particularly relevant in modelling biological transport and reaction-diffusion processes. They describe the movement of substances or entities within biological systems and the interactions between different components. Based on the mass conservation law, many mathematical models have been studied in cellular populations buttenschon2018 . For example, if $u(\mathbf{x},t)$ and $J(\mathbf{x},t)$ are the density and flux of a population, then the conservation equation is given by
@@
An appropriate choice of flux helps in capturing the different biological phenomena. As an example, the flux $J$ can be divided into the random motion ($J_{d}$) and the adhesion ($J_{a}$), i.e.,
@@
We assume Fick’s law for the diffusive flux, i.e., $J_{d}=-d_{1}\nabla u$. On the other hand, following Armstrong et al. armstrong2006 , the adhesion flux is proportional to the population density and adhesion force. It is inversely proportional to the cell size, so
@@
${}u_{t}(\mathbf{x},t)=d_{1}\Delta u(\mathbf{x},t)-\nabla\cdot(u(\mathbf{x},t)(\Phi\ast u)(\mathbf{x},t)),$ (20)
@@
In sherratt2009 , authors have considered the nonlocal model (20) with an additional term (logistic growth) in the reaction kinetics. They derived sufficient conditions for the boundedness of the solutions, but these conditions do not satisfy numerically for some parameter values, and hence, there are still some mathematical challenges.
@@ 2.11 Nonlocal variational problems
In traditional variational problems, the goal is to find an extremum of a function that depends on local derivatives of a function. Nonlocal variational problems are mathematical optimization problems that involve nonlocal operators in their formulations. These formulations have been used to derive and study nonlocal PDEs in diverse physical and biological systems. The general form of a nonlocal variational problem can be expressed as follows:
@@
Nonlocal mathematical model also arises from crop-raiding of large-bodied mammals living in the biodiversity-rich tropics hsu2014 . Let $\Omega$ be a spatial domain with $\overline{\Omega}_{0}\subset\Omega$ and $u(\mathbf{x},t)$ be the population density of the mammal at position $\mathbf{x}$ and time $t$. Here, we assume that the forest is safe for the mammal species but is of poor resources, while the farm has rich resources but is dangerous. Furthermore, mammals cannot survive if they only stay in the forest and do not attempt to go out to search for food; they cannot produce offspring on the farm because it is not a safe place for reproduction. In this case, the model takes the form hsu2014 :
@@
The parameter $d$ is the diffusivity, and $p\geq 2$ takes into account the over-crowding effects of the population in $\Omega$. The non-negative steady-states for the model (22) are the nonnegative solutions of the semilinear nonlocal elliptic problem
@@
A similar type of Kirchhoff-Carrier type equation boundary value problems studied in the literature xu2022 :
@@
where $|u|_{\gamma}=(\int_{\Omega}|u|^{\gamma})^{1/\gamma}$, $||u||=(\int_{\Omega}|\nabla u|^{2})^{1/2}$, $\alpha$ is a non-negative function, and $c$ is a non-negative continuous function in $\Omega$. The conditions on $f$ and the other parameters are given in xu2022 . A weak solution to this problem is a function $u$ in a Hilbert space $H_{0}^{1}(\Omega)$ such that
@@
Structure-preserving characteristics and variational integrators are fundamental notions in numerical analysis, particularly in the simulation of dynamical systems leimkuhler2004 . These approaches are intended to retain fundamental aspects of the underlying physical system, such as energy conservation, ensuring that the numerical solution closely resembles the genuine system’s behaviour over long time periods. These methodologies are used in a variety of scientific areas, ranging from celestial mechanics to molecular dynamics simulations.
Sometimes, the complete information of a system’s states may be inaccessible, but only partial information about the states is observable due to limitations in experimental techniques. In this case, observability is a critical aspect of partially observed dynamics, highlighting which parts of a system are accessible to measurement luenberger1971 . The reduction techniques focus on accounting for hidden variables and incorporating partial observations into the system’s modelling. Stochastic models and Bayesian inference are employed to handle uncertainties associated with partially observed dynamics, and these frameworks provide probabilistic representations of hidden states and update beliefs based on observed data.
@@ 2.12 Nonlocal models in pattern formations
Pattern analysis was essential to human survival and advancement in the past. In order to make sense of the world, ancient civilizations studied patterns in astronomy, agriculture, social behaviour, and nature. Seasonal trends and natural cycles were intimately related to agricultural operations. To keep the soil fertile and minimize depletion, farmers would rotate the crops they grew in a given field. Numerous ancient societies, like the Greeks, Egyptians, and Mesopotamians, depended on the yearly inundation of rivers like the Nile and Tigris-Euphrates to replenish the soil with nutrients and used astronomical patterns as a guide. Although the first agricultural communities probably lacked the sophisticated maps we have today, they did manage and plan their agricultural activities using simple spatial organizing techniques. For instance, the historic Indus Valley civilization metropolis of Mohenjo-Daro which dates to approximately 2600 BCE. Cuneiform tablets uncovered in ancient Mesopotamia, one of the first centres of civilization, date to approximately 2300 BCE and show primitive maps of land property.
@@
History demonstrates that the time of agricultural activities was greatly influenced by the locations of the sun, moon, and stars. A Greek mathematician and astronomer, Hipparchus (190–120 BCE), is frequently acknowledged for having made some of the first attempts at star mapping. By mapping the stars using a coordinate system, he popularized the ideas of celestial longitude and latitude. Hipparchus’s suggestion of a method to measure the apparent brightness of stars is one of his most important contributions. Babylonian mathematicians also investigated natural numerical patterns. Pythagorean triples are listed on the Plimpton 322 clay tablet from ancient Babylonia, which dates to approximately 1800 BCE and illustrates mathematical patterns in ancient times. In addition, traditional medical professionals and healers examined trends in disease and symptomology. They helped to shape the early forms of herbal medicine and other medical practices by observing the impact of specific plants or ceremonies on health. Ancient civilizations didn’t have the advanced technology that we use today, but they did develop their own methodologies to understand the world. Their models may not have been explicitly nonlocal in the way we think about it today, but they certainly reflected an understanding of interconnectedness and patterns in the world.
@@
Nonlocal interactions in ecological systems can have a significant impact on spatiotemporal pattern formation, e.g., the development of spatial patterns in plants. Nonlocal interactions between plants, such as resource competition or facilitation, can result in the creation of regular patterns such as stripes, spots, or waves. These patterns frequently emerge as a result of self-organization in which interactions between individuals on a broader scale give rise to organized patterns on a landscape level. Nonlocal interactions in predator-prey relationships can result in the development of predator fronts, in which predators pursue prey concentrations over longer distances, altering the spatial dynamics of both populations. The nonlocal predator-prey interactions are generally two types: intra- and inter-specific. First, we consider a general nonlocal prey-predator model with nonlocal interaction in the intra-specific prey competition in a two-species population as
@@
subjected to non-negative initial conditions and appropriate boundary conditions of prey ($u$) and predator ($v$) populations. The parameters $\alpha$ and $\kappa$ are intrinsic growth rate and environmental carrying capacity, respectively, and $e$ $(0<e\leq 1)$ is the conversion efficiency from prey biomass into predator biomass. If $e=0$, then there is no interaction between the species. For simplicity, we have considered a one-dimensional spatial domain $\Omega\subset\mathbb{R}$, and $d_{1}$ and $d_{2}$ are the self-diffusion coefficients. As defined earlier, $\phi\ast u$ is the convolution term. The kernel function $\phi$ satisfies all the assumptions mentioned earlier, i.e., normalized, and has compact support in $\mathbb{R}$. For an unbiased movement of the species in both spatial directions, we assume $\phi$ as an even function. Still, sometimes, species do not have the same area for nonlocal consumption of resources in both spatial directions. This generally happens for a bounded spatial domain, e.g., studying a nonlocal interaction for the populations living in a river sherratt2016 ; sun2019 ; lutscher2005 . The function $f(u)$ is called a functional response, and it satisfies three conditions myerscough1996 ; pal2021cnsns : (i) $f(0)=0$, (ii) $f$ is an increasing function and (iii) there exists a finite $M>0$ such that $\lim_{u\rightarrow\infty}f(u)=M$. Many biologically relevant functional responses are available in the literature, e.g., Holling type II and III, Ivlev function, etc. myerscough1996 ; pal2021cnsns ; real1977 ; skalski2001 . The function $g(v)$ represents the per capita death rate of the predator population berryman1992 . In banerjee2017ec , authors have studied a particular form of the nonlocal model (23). Nevertheless, the local model corresponding to that nonlocal model does not produce any nonhomogeneous stationary pattern, but the nonlocal model does.
Researchers have also studied the pattern formation for nonlocal models with more than two species populations. Generally, the three components of the local or nonlocal model in ecology add another feature to the two-species model. It allows us to study the dynamics of the prey-predator model in the presence of an additional species. These three components of the model may be (a) food chain; (b) two predators–one prey; (c) one predator–two prey; (d) food chain with omnivory; and (e) food chain with cycle hsu2015 . Autry et al. autry2018 have considered a three-species nonlocal RD model in a food chain system with ratio-dependent functional responses as:
@@
where $u$, $v$, and $w$ represent the concentration of prey, predator, and super-predator, respectively. All the convolution terms are defined as before, and $\phi$ and $\psi$ are the kernel functions. Here, the prey and super-predator interact directly with the intermediate predator and only interact indirectly with the intermediate predator. Authors in autry2018 have studied two aspects of the nonlocality: C-type and P-type. The $C$-type nonlocality occurs when the prey species competes with itself. This type of nonlocal interaction usually occurs due to the overlapping ranges of the species, e.g., extensive root systems or a shared water table in plant species. On the other hand, P-type stands for pest type, which arises when one species nonlocally interacts with the other species, and this nonlocal interaction only affects the encounter rate, which appears in the functional response term. In addition, one can consider the same type of nonlocal term in the functional response for the first and second equations to study the nonlocal effect of high-mobile intermediate predator species $v$ on the crops or prey $u$.
In population biology, age structure models and pattern formation provide insights into the dynamics of populations and their structural changes over time. Age structure refers to the distribution of individuals in a population across different age groups, and understanding its patterns is crucial for demographic analysis and prediction. McKendrick and Von Foerster developed this concept m1925 ; von1959 . Suppose $n(a,t)$ is the population at age $a$ and time $t$. Following McKendrick - Von Foerster, the population $n(a,t)$ satisfies the following equation trucco1965 :

### v2

@@ 2.7 Nonlocal epidemiological models and nonlocality in immunology
Nonlocal epidemiological models have been used to study the spread of infectious diseases in biological systems, considering interactions beyond immediate neighbours and capturing the effects of long-range dispersal and spatial dependencies. In 1927, Kermack and McKendrick first proposed a deterministic SIR model to capture a transmitted viral or bacterial agent in a closed population [15, 134]. Let $S(t)$ be the density of a population susceptible to a disease but not yet infected at the time $t\geq 0$. Suppose $A(\theta)$ is the expected infectivity of an individual infected $\theta$ time ago. Then, the Kermack-McKendrick model follows the integral differential equation [15]:
@@
describes the epidemic’s growth in the initial phase of the disease propagation. The invasion criteria of the epidemic is $R_{0}$, which is greater than one. For the special choice $A(\theta)=\beta\exp(-\gamma\theta)$, where $\beta,\gamma>0$, and the quantity
@@
$\frac{dR}{dt}$ $=\gamma I(t),$
with non-negative initial conditions. This dynamical system provides changes in density among susceptible, infected and recovered compartments over time. Here, the individuals can be treated as the particles in a system where their interactions lead to collective behaviour, e.g., the spread of disease. In 1965, Kendall extended this SIR model (11) into a nonlocal epidemic model [57]. Let $S(x,t)$, $I(x,t)$, and $R(x,t)$ be the densities of the susceptible, infected and removed individuals at the spatial location $x\in\mathbb{R}$ and time $t\geq 0$ such that their sum $S+I+R$ remains time-independent. We assume that the infected individuals are infectious and that the rate of infection is
@@
where $\beta$ is a positive constant and $\Psi$ is the normalized kernel function in $\mathbb{R}$. This mechanism considers that the infection rate at a location is influenced by contributions from distant regions, and the kernel defines the interaction radius or probability of long-range infection spread. This function takes various forms, such as exponential decay, where the influence of an infected individual decreases with distance. Taking these into the SIR model, the modified model becomes
@@
In classical RD models, the infections spread locally and gradually to adjacent regions. However, the nonlocal interaction leads to wave-like infection propagation due to the fact that distant regions can become infected simultaneously or even ahead of nearby regions, depending on the strength and range of nonlocal interactions (e.g., travel networks).
Nonlocal epidemiological models have been applied in biology to study the spread of infectious diseases in spatially extended and heterogeneous environments, such as urban areas with varying population densities or regions with different levels of healthcare access [44, 135, 136, 137]. These models incorporate long-range interactions and spatial dependencies, allowing for a more realistic representation of disease transmission dynamics. The COVID-19 pandemic was a global health crisis caused by the SARS-CoV-2 virus in a heterogeneous environment [138]. This disease spreads rapidly through respiratory droplets when an infected person talks, coughs, or sneezes, and new variants of the virus continuously emerge, which causes concern among health experts worldwide. Different types of measurements have been implemented to control the transmission of the virus, including targeted lockdowns, ramped-up testing, and widespread vaccination efforts. In terms of mathematical modelling, numerous mathematical frameworks have been proposed to investigate the transmission dynamics of the COVID-19 pandemic [139, 140, 141], and we highlight one of them where the spatial domain is considered. The novel coronavirus infects contacts during cold-chain transmission, so studying the characteristics of COVID-19 dynamics can be more accurate in considering long-range population diffusion. Suppose $S(x,t)$, $V(x,t)$, $E(x,t)$, $I_{a}(x,t)$, and $I_{s}(x,t)$ are respectively for the populations of susceptible, fully vaccinated, latent infection, asymptomatic infected, and symptomatic infected at a point $x\in\mathbb{R}$ and time $t>0$, with $B(x,t)$ as the SARS-CoV-2 virus in the environment. Following [140], we consider the spatio-temporal nonlocal mathematical model for COVID-19 as:
@@
where $\psi(x-y)$ is the probability that an individual at position $y$ moves to position $x$ in the space $\Omega$. This type of nonlocal diffusion accounts for jump-like interactions, where infection can jump across large distances (such as through transportation networks or migration). The functions $f_{s}(x,t)$ and $f_{v}(x,t)$ are defined as
@@
where $\beta_{k}(k=e,a,s)$ are the rate of infection of infected people to susceptible people and $\alpha_{k}(k=e,a,s)$ are the rate of infection of infected people to those vaccinated, and $\beta_{b}$ and $\alpha_{b}$ are the rate of infection of the virus in the population. The other parameters used in the model (13) are given in [140]. The authors have shown that the effective vaccination rate is one of the best measures to prevent and control the transmission of COVID-19 under current medical conditions.
Nonlocality in immunology refers to the phenomena in which immune system interactions and signalling processes extend beyond the immediate area of the immune cells involved. T-cell receptor (TCR) activation, for example, refers to the assumption that T-cell responses are not limited to the local contact between the TCR and its antigen. T-cell activation, on the other hand, is a networked and dynamic process that includes interactions at remote places [142]. It can lead to systemic effects, influencing immune responses and inflammation throughout the body, which has implications for autoimmune diseases and chronic inflammatory conditions [143]. Furthermore, nonlocal TCR activation is important in cancer immunotherapy because activated T cells may target cancer cells in distant metastatic locations. [144]. In T memory cell homeostasis, for example, the bone marrow serves as a niche and interacts with stromal cells, and cytokines contribute to their survival and nonlocal homeostasis [145]. These T cells grow and contract in distinct stages, including interactions and signals that reach beyond the T cells’ local surroundings. During T cell activation and proliferation, cytokines such as interleukins are generated not just locally at the site of activation but also systemically. These cytokines assist in the nonlocal regulation of T-cell proliferation [146].
@@ 2.8 Relaxation-time models in life sciences
Relaxation-time models are pretty interesting, especially in the context of life sciences. These models help us understand the dynamics of biological systems and provide insights into the time-dependent behaviour of various processes. For instance, the nonlocal model in space and time can be considered the relaxation-time nonlocal model (5) in space and time. A similar type of extension can be applied to the model (12). We first include the spatial variable into the model (12) to account for the random movement of individuals as [147]:
@@
where the kernel function describes the interaction between the infective and the susceptible individuals at location $x$ and the present $t$, which occurred at location $y$ and at earlier instance $s$. This kernel function satisfies some conditions mentioned in [134]. In this context, the term “relaxation” refers to the process by which the population returns to equilibrium after a disruption, such as an epidemic outbreak. In a nonlocal SIR model, the relaxation process includes both local recovery rates and nonlocal temporal impacts. For example, after a local epidemic has subsided, a population may continue to be infected because adjacent regions continue to have high infection rates, delaying the return to equilibrium. In contrast, a local model would predict a faster return to equilibrium once the local infection subsides.
The nonlocal effects and time delay may come in the model when the mature death and diffusion rates are age-independent. To derive such a delay model, we follow the work [148] where the author considered a single-species population model with two age classes and a fixed maturation period living in a spatial transport field. Suppose $u(x,a,t)$ is the population density at a spatial point $x\in\Omega\subset\mathbb{R}$ and time $t\geq 0$ with age $a\geq 0$. Following [148], the single-species model given by
@@
where $v^{s}(x,t)=u(x,t-s,t)$. The dynamics of a population can be heavily influenced by age-specific behaviours [149, 150, 151]. For instance, younger individuals may disperse more than older ones, older individuals often have higher mortality rates, etc.
@@ 2.9 Nonlocal problems and conservation laws
Nonlocal problems and conservation laws are particularly relevant in modelling biological transport and reaction-diffusion processes. They describe the movement of substances or entities within biological systems and the interactions between different components. Traditional conservation laws (e.g., conservation of mass, momentum, energy) express that certain quantities remain constant within a closed system. They are typically local, depending only on the properties and states of nearby points. Based on the mass conservation law, many mathematical models have been studied in cellular populations [152, 153]. For example, if $u(\mathbf{x},t)$ and $J(\mathbf{x},t)$ are the density and flux of a population, then the conservation equation is given by
@@
An appropriate choice of flux helps capture different biological phenomena. As an example, the flux $J$ can be divided into the random motion ($J_{d}$) and the adhesion ($J_{a}$), i.e.,
@@
We assume Fick’s law for the diffusive flux, i.e., $J_{d}=-d\nabla u$. On the other hand, following Armstrong et al. [154], the adhesion flux is proportional to the population density and adhesion force. It is inversely proportional to the cell size, so
@@
${}u_{t}(\mathbf{x},t)=d\Delta u(\mathbf{x},t)-\nabla\cdot(u(\mathbf{x},t)(\Phi\ast u)(\mathbf{x},t)),$ (18)
@@
Analyzing the stability and well-posedness of solutions to nonlocal conservation laws is more complex than local laws, often requiring specialized techniques to ensure unique and stable solutions [155, 156, 157]. The authors in [155] considered the nonlocal model (18) with an additional term (logistic growth) in the reaction kinetics and derived sufficient conditions for the boundedness of the solutions. However, these conditions do not satisfy some parameter values numerically, and hence, there are still some mathematical challenges.
@@ 2.10 Nonlocal variational problems
In traditional variational problems, the goal is to find an extremum of a function that depends on local derivatives of a function. Nonlocal variational problems are mathematical optimization problems that involve nonlocal operators in their formulations. These formulations have been used to derive and study nonlocal partial differential equations (PDEs) in diverse physical and biological systems. The general form of a nonlocal variational problem can be expressed as follows:
@@
Nonlocal mathematical model also arises from crop-raiding of large-bodied mammals living in the biodiversity-rich tropics [158]. Let $\Omega$ be a spatial domain with $\overline{\Omega}_{0}\subset\Omega$ and $u(\mathbf{x},t)$ be the population density of the mammal at position $\mathbf{x}$ and time $t$. Here, we assume that the forest is safe for the mammal species but is of poor resources, while the farm has rich resources but is dangerous. Furthermore, mammals cannot survive if they only stay in the forest and do not attempt to go out to search for food; they cannot produce offspring on the farm because it is not a safe place for reproduction. In this case, the model takes the form [158]:
@@
The parameter $d$ is the diffusivity and $p\geq 2$ considers the population’s overcrowding effects in $\Omega$. The non-negative steady-states for the model (20) are the nonnegative solutions of the semilinear nonlocal elliptic problem
@@
A similar type of Kirchhoff-Carrier type equation boundary value problems studied in the literature [159]:
@@
where $|u|_{\gamma}=(\int_{\Omega}|u|^{\gamma})^{1/\gamma}$, $||u||=(\int_{\Omega}|\nabla u|^{2})^{1/2}$, $\alpha$ is a non-negative function, and $c$ is a non-negative continuous function in $\Omega$. The conditions on $f$ and the other parameters are given in [159]. A weak solution to this problem is a function $u$ in a Hilbert space $H_{0}^{1}(\Omega)$ such that
@@
Structure-preserving characteristics and variational integrators are fundamental notions in numerical analysis, particularly in the simulation of dynamical systems [160]. These approaches are intended to retain fundamental aspects of the underlying physical system, such as energy conservation, ensuring that the numerical solution closely resembles the genuine system’s behaviour over long periods. These methodologies are used in various scientific areas, ranging from celestial mechanics to molecular dynamics simulations.
Sometimes, the complete information of a system’s states may be inaccessible, but only partial information about the states is observable due to limitations in experimental techniques. In this case, observability is a critical aspect of partially observed dynamics, highlighting which parts of a system are accessible to measurement [161]. The reduction techniques focus on accounting for hidden variables and incorporating partial observations into the system’s modelling. Stochastic models and Bayesian inference are employed to handle uncertainties associated with partially observed dynamics, and these frameworks provide probabilistic representations of hidden states and update beliefs based on observed data.
@@ 2.11 Nonlocal models in pattern formations
Pattern analysis was essential to human survival and advancement in the past. To make sense of the world, ancient civilizations studied patterns in astronomy, agriculture, social behaviour, and nature. Seasonal trends and natural cycles were intimately related to agricultural operations. Farmers rotated the crops to maintain soil fertility and minimize depletion, and they grew in a given field. Numerous ancient societies, like the Greeks, Egyptians, and Mesopotamians, depended on the yearly inundation of rivers like the Nile and Tigris-Euphrates to replenish the soil with nutrients and used astronomical patterns as a guide. The early agricultural communities may not have had the advanced maps that we have today, but they still effectively planned and organized their agricultural activities using basic spatial techniques. For example, the ancient city of Mohenjo-Daro dates back to around 2600 BCE and was part of the historic Indus Valley civilization. Cuneiform tablets uncovered in ancient Mesopotamia, one of the first centres of civilization, date to approximately 2300 BCE and show primitive maps of land property.
@@
History demonstrates that the time of agricultural activities was greatly influenced by the locations of the sun, moon, and stars. A Greek mathematician and astronomer, Hipparchus (190–120 BCE), is frequently acknowledged for having made some of the first attempts at star mapping. By mapping the stars using a coordinate system, he popularized the ideas of celestial longitude and latitude. Hipparchus’s suggestion of a method to measure the apparent brightness of stars is one of his most important contributions. Babylonian mathematicians also investigated natural numerical patterns. Pythagorean triples are listed on the Plimpton 322 clay tablet from ancient Babylonia, which dates to approximately 1800 BCE and illustrates mathematical patterns in ancient times. In addition, traditional medical professionals and healers examined trends in disease and symptomology. They helped to shape the early forms of herbal medicine and other medical practices by observing the impact of specific plants or ceremonies on health. Ancient civilizations did not have the advanced technology that we use today, but they developed their methodologies to understand the world. Their models may not have been explicitly nonlocal in the way we think about it today, but they certainly reflected an understanding of interconnectedness and patterns in the world.
@@
The physics of ecological pattern formation explores the complex dynamics between consumer-resource interactions, which often leads to various spatial and temporal structures [162]. Nonlocal interactions in ecological systems can significantly impact such spatio-temporal pattern formations, e.g., the development of spatial patterns in plants. Nonlocal interactions between plants, such as resource competition or facilitation, can result in the creation of regular patterns such as stripes, spots, or waves. These patterns frequently emerge as a result of self-organization in which interactions between individuals on a broader scale give rise to organized patterns on a landscape level. Nonlocal interactions in predator-prey relationships can result in the development of predator fronts, in which predators pursue prey concentrations over longer distances, altering the spatial dynamics of both populations. Nonlocal predator-prey interactions are generally two types: intra- and inter-specific. First, we consider a general nonlocal predator-prey model with nonlocal interaction in the intra-specific prey competition in a two-species population as
@@
subjected to non-negative initial conditions and appropriate boundary conditions of prey ($u$) and predator ($v$) populations. The parameters $\alpha$ and $\kappa$ are intrinsic growth rate and environmental carrying capacity, respectively, and $e$ $(0<e\leq 1)$ is the conversion efficiency from prey biomass into predator biomass. If $e=0$, then there is no interaction between the species. For simplicity, we have considered a one-dimensional spatial domain $\Omega\subset\mathbb{R}$, and $d_{1}$ and $d_{2}$ are the self-diffusion coefficients. As defined earlier, $\phi\ast u$ is the convolution term. The kernel function $\phi$ satisfies all the assumptions mentioned earlier, i.e., normalized, and has compact support in $\mathbb{R}$. For an unbiased movement of the species in both spatial directions, we assume $\phi$ as an even function. Still, sometimes, species do not have the same area for nonlocal consumption of resources in both spatial directions. This generally happens for a bounded spatial domain, e.g., studying a nonlocal interaction for the populations living in a river [54, 163, 164]. The function $f(u)$ is called a functional response, and it satisfies three conditions [165, 166]: (i) $f(0)=0$, (ii) $f$ is an increasing function and (iii) there exists a finite $M>0$ such that $\lim_{u\rightarrow\infty}f(u)=M$. Many biologically relevant functional responses are available in the literature, e.g., Holling type II and III, Ivlev function, etc. [165, 166, 167, 168, 169, 170, 51, 171]. The function $g(v)$ represents the per capita death rate of the predator population [172]. In [90], authors have studied a particular form of the nonlocal model (21). Nevertheless, the local model corresponding to that nonlocal model does not produce any nonhomogeneous stationary pattern, but the nonlocal model does.
Researchers have also studied the pattern formation for nonlocal models with more than two species populations. Generally, the three components of the local or nonlocal model in ecology add another feature to the two-component model [173]. It allows us to study the dynamics of the predator-prey model in the presence of an additional species. These three components of the model may be (a) food chain; (b) two predators–one prey; (c) one predator–two prey; (d) food chain with omnivory; and (e) food chain with cycle [174]. Autry et al. [175] have considered a three-species nonlocal RD model in a food chain system with ratio-dependent functional responses as:
@@
where $u$, $v$, and $w$ represent the concentration of prey, predator, and super-predator, respectively. All the convolution terms are defined as before, and $\phi$ and $\psi$ are the kernel functions. Here, the prey and super-predator interact directly with the intermediate predator and only interact indirectly with the intermediate predator. Authors in [175] have studied two aspects of the nonlocality: C-type and P-type. The C-type nonlocality occurs when the prey species competes with itself. This type of nonlocal interaction usually occurs due to the overlapping ranges of the species, e.g., extensive root systems or a shared water table in plant species. On the other hand, P-type stands for pest type, which arises when one species nonlocally interacts with the other species, and this nonlocal interaction only affects the encounter rate, which appears in the functional response term. In addition, one can consider the same type of nonlocal term in the functional response for the first and second equations to study the nonlocal effect of high-mobile intermediate predator species $v$ on the crops or prey $u$.
In population biology, age structure models and pattern formation provide insights into the dynamics of populations and their structural changes over time. Age structure refers to the distribution of individuals in a population across different age groups, and understanding its patterns is crucial for demographic analysis and prediction. McKendrick and Von Foerster developed this concept [176, 177]. Suppose $n(a,t)$ is the population at age $a$ and time $t$. Following McKendrick-Von Foerster, the population $n(a,t)$ satisfies the following equation [178]: